AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 397     Entry time: Thu May 11 21:22:54 2023
Author: RSS, TD 
Subject: V-I test 

Today, the foil was opened, the pins on the bottom right detector were checked and looked fine, the right-hand arm was moved a bit in and then out, and the foil was tightly closed again. The negative bias voltage was applied from 0 to 50 V in steps of 10 V. During the measurement, the lights in the cave were switched off to ensure no leakage current from the light*. The results are shown in Attachment 1. To condition the detector, it was left at a negative bias voltage of 50 V for ~ 3 hours with lights on. No increase in current was seen (see Attachment 2), indicating the chamber was nicely covered with foil. Then the negative voltage was increased from 60 V to 150 V in steps of 10 V. No large current readings, like yesterday, were seen today (see Attachment 3). The detector was then left at -150 V for ~ 30 minutes with the current to be nearly stable with fluctuations in the range of 70-100 nA. The foil was then taken off to check if there are any dust particles on the detector surface. As can be seen in Attachment 6, some dust particle was found at the bias/multi-guard ring (MGR) boundary structures of the DSSSD. The area was then nicely cleaned with isopropanol + lab tissue wipe (see Attachment 7) and the chamber was covered with aluminum foil again. The negative bias voltage was applied from 0 to 150 V in steps of 10 V again and the results can be seen in attachment 5, indicating the response of the detector to be normal, unlike yesterday. Before calling the day off, the detector was left at -150 V overnight.

 

* Every time, before switching off the lights, it was checked that no one else was in the cave.

Attachment 1: test1.png  17 kB  Uploaded Thu May 11 22:23:02 2023  | Hide | Hide all
test1.png
Attachment 2: test2.png  22 kB  Uploaded Thu May 11 22:23:10 2023  | Hide | Hide all
test2.png
Attachment 3: test3.png  18 kB  Uploaded Thu May 11 22:23:19 2023  | Hide | Hide all
test3.png
Attachment 4: test4.png  22 kB  Uploaded Thu May 11 22:23:27 2023  | Hide | Hide all
test4.png
Attachment 5: test5.png  18 kB  Uploaded Thu May 11 22:24:47 2023  | Hide | Hide all
test5.png
Attachment 6: IMG_8800.jpg  349 kB  Uploaded Thu May 11 22:37:23 2023  | Hide | Hide all
IMG_8800.jpg
Attachment 7: IMG_8804.jpg  986 kB  Uploaded Thu May 11 22:37:36 2023  | Hide | Hide all
IMG_8804.jpg
ELOG V3.1.4-unknown